# Ultra-Slow-Roll Inflation on the Lattice: Backreaction and Nonlinear Effects

### v1

@@ Abstract
Violating the slow-roll regime during the final stages of inflation can significantly enhance curvature perturbations, a scenario often invoked in models producing primordial black holes and small-scale scalar induced gravitational waves. When perturbations are enhanced, one approaches the regime in which tree-level computations are insufficient, and nonlinear corrections may become relevant. In this work, we conduct lattice simulations of ultra-slow-roll (USR) dynamics to investigate the significance of nonlinear effects, both in terms of backreaction on the background and in the evolution of perturbations. Our systematic study of various USR potentials reveals that nonlinear corrections are significant when the tree-level curvature power spectrum peaks at $\mathcal{P}^{\rm max}_{\zeta}={\cal O}(10^{-3})-{\cal O}(10^{-2})$, with 5%--10% corrections. Larger enhancements yield even greater differences. We establish a universal relation between simulation and tree-level quantities $\dot{\phi}=\dot{\phi}_{\rm tree}\left(1+\sqrt{\mathcal{P}^{\rm max}_{\zeta,\rm tree}}\right)$ at the end of the USR phase, which is valid in all cases we consider. Additionally, we explore how nonlinear interactions during the USR phase affect the clustering and non-Gaussianity of scalar fluctuations, crucial for understanding the phenomenological consequences of USR, such as scalar-induced gravitational waves and primordial black holes. Our findings demonstrate the necessity of going beyond leading order perturbation theory results, through higher-order or non-perturbative computations, to make robust predictions for inflation models exhibiting a USR phase.
@@ II.3.2 Action for perturbations with and without Wands duality
Note that in the decoupling limit that we consider, one simply has $V_{2}/H^{2}\simeq m^{2}_{\textrm{eff}}/H^{2}\simeq 9/4-\nu^{2}$. Higher-order derivatives simply follow from $V_{n+1}=\dot{\phi}V_{n}$. [^4] Notice that Ballesteros and Egea (2024) implicitly assumed a positive inflaton velocity and hence does not have $\textrm{sign}(\dot{\phi})$ in the corresponding Eq. (11b).
@@ III.2.1 Background dynamics
First, we compare the background evolution with the solution of the homogeneous Klein-Gordon equation (4) that we obtain numerically. This comparison is shown in the top panels of Figs. 5, 6, and 7 for each of the three cases. In all cases, the background solution agrees with SPT at the $<1\%$ level for $\mathcal{P}^{\rm max}_{\zeta,\rm tree}=10^{-4}$. However, for larger values of $\mathcal{P}^{\rm max}_{\zeta,\rm tree}$, the inflaton mean velocity $\langle\dot{\phi}\rangle$ in the lattice simulation deviates from SPT predictions, with O(1) deviations for $\mathcal{P}^{\rm max}_{\zeta,\rm tree}=1$. Moreover, during the USR phase, the velocity in the simulations does not decrease as much as predicted by the purely homogeneous evolution. This is a consequence of the backreaction of fluctuations on the background evolution, which is not taken into account by SPT. In Sec. III.2.3, we give tentative explanations of this phenomenon. Due to the very small inflaton velocity throughout the USR phase, the backreaction effect in the mean field value $\langle\phi\rangle$ as a function of number of $e$-folds is much less significant, but the two are naturally related, as the phase-space portrait in Fig. 10 will show explicitly. In particular, the field evolution returns to the SR attractor after the USR phase.
@@ IV Conclusions and outlook
It was recently suggested that USR dynamics advocated by some of the PBH formation scenarios may violate perturbativity and induce loop corrections affecting much longer modes associated to the scales observed through the CMB Kristiano and Yokoyama (2024c). [^7] Instead, see Fumagalli et al. (2024) for effects at near infrared scales when the power spectrum enhancement is due to particle production. While the existence and magnitude of this effect is still under debate Riotto (2023a); Kristiano and Yokoyama (2024d); Riotto (2023b); Kristiano and Yokoyama (2024b); Franciolini et al. (2024b); Davies et al. (2024); Firouzjahi and Riotto (2024); Iacconi et al. (2024); Motohashi and Tada (2023); Tasinato (2024, 2023); Firouzjahi (2023); Fumagalli (2023); Tada et al. (2024); Firouzjahi (2024); Braglia and Pinol (2024); Kawaguchi et al. (2024a); Ballesteros and Egea (2024); Inomata (2024); Fumagalli (2024); Kawaguchi et al. (2024b); Green and Gupta (2024), we are not able to address this question due to the relatively small range of scales that can be tracked in the lattice box. However, we are in the position to comment on some of the related issues. In particular, our result shows perturbativity is still not completely violated as long as the amplitude of perturbations are smaller than order unity (contrary to what happens in oscillatory potentials, where perturbativity is completely broken already for $\mathcal{P}_{\zeta,\rm tree}^{\rm max}=10^{-2}$ Caravano et al. (2024)). Also, we confirm the existence of important backreaction on the background dynamics, typically captured by tadpole contributions, which were suggested to play a crucial role in protecting long modes from loop corrections Inomata (2024); Fumagalli (2024).
@@
Looking forward, it would be interesting to extend the present work along different directions. Firstly, it will be important to derive the full non-Gaussian statistics for the curvature perturbation, including the nonlinear relation between $\zeta$ and $\delta\phi$ from our lattice simulations. Other groups attempted to study the effect of backreaction and nonlinearities in the USR scenarios employing full in-in computations or the stochastic inflation formalism (see e.g. Pattison et al. (2017); Biagetti et al. (2018); Ezquiaga et al. (2020); Ballesteros et al. (2020); Pattison et al. (2021); Biagetti et al. (2021); Figueroa et al. (2022); Raatikainen et al. (2024); Mizuguchi et al. (2024); Ballesteros and Egea (2024); Animali and Vennin (2024); Jackson et al. (2024)), and it would be interesting to compare these techniques with lattice results. Additionally, evaluating the four-point function of the field fluctuations $\delta\phi$ would provide key insights into the non-perturbative source of SIGW (complementing recent efforts in this direction Cai et al. (2019); Unal (2019); Yuan and Huang (2021); Atal and Domènech (2021); Adshead et al. (2021); Abe et al. (2023); Chang et al. (2023); Garcia-Saenz et al. (2023); Li et al. (2023); Perna et al. (2024); Ruiz and Rey (2024)). These next steps will allow us to better quantify the importance of nonlinearities in the USR scenarios, paving the way for robust observational predictions for GW experiments, as well as constraints on the PBH abundance.
@@ Acknowledgements.
We thank Guillermo Ballesteros, Keisuke Inomata, Yuichiro Tada and Vincent Vennin for insightful discussions. A.C. acknowledges funding support from the Initiative Physique des Infinis (IPI), a research training program of the Idex SUPER at Sorbonne Université. During the course of this work, S.RP was supported by the European Research Council under the European Union’s Horizon 2020 research and innovation programme (grant agreement No 758792, Starting Grant project GEODESI). This article is distributed under the Creative Commons Attribution International Licence (CC-BY 4.0).

### v2

@@ Abstract
Violating the slow-roll regime during the final stages of inflation can significantly enhance curvature perturbations, a scenario often invoked in models producing primordial black holes and small-scale scalar induced gravitational waves. When perturbations are enhanced, one approaches the regime in which tree-level computations are insufficient, and nonlinear corrections may become relevant. In this work, we conduct lattice simulations of ultra-slow-roll (USR) dynamics to investigate the significance of nonlinear effects, both in terms of backreaction on the background and in the evolution of perturbations. Our systematic study of various USR potentials reveals that nonlinear corrections are significant when the tree-level curvature power spectrum peaks at $\mathcal{P}^{\rm max}_{\zeta}={\cal O}(10^{-3})-{\cal O}(10^{-2})$, with 5%--10% corrections. Larger enhancements yield even greater differences. We find a simple universal relation between simulation and tree-level quantities $\dot{\phi}=\dot{\phi}_{\rm tree}\left(1+\sqrt{\mathcal{P}^{\rm max}_{\zeta,\rm tree}}\right)$ at the end of the USR phase, which is valid in all cases we consider. Additionally, we explore how nonlinear interactions during the USR phase affect the clustering and non-Gaussianity of scalar fluctuations, crucial for understanding the phenomenological consequences of USR, such as scalar-induced gravitational waves and primordial black holes. Our findings demonstrate the necessity of going beyond leading order perturbation theory results, through higher-order or non-perturbative computations, to make robust predictions for inflation models exhibiting a USR phase.
@@ II.3.2 Action for perturbations with and without Wands duality
Note that in the decoupling limit that we consider, one simply has $V_{2}/H^{2}\simeq m^{2}_{\textrm{eff}}/H^{2}\simeq 9/4-\nu^{2}$. Higher-order derivatives simply follow from $\dot{V}_{n}=\dot{\phi}V_{n+1}$. [^4] Notice that Ballesteros and Egea (2024) implicitly assumed a positive inflaton velocity and hence does not have $\textrm{sign}(\dot{\phi})$ in the corresponding Eq. (11b).
@@ III.2.1 Background dynamics
First, we compare the background evolution with the solution of the homogeneous Klein-Gordon equation (4) that we obtain numerically. This comparison is shown in the top panels of Figs. 5, 6, and 7 for each of the three cases. In all cases, the background solution agrees with SPT at the $<1\%$ level for $\mathcal{P}^{\rm max}_{\zeta,\rm tree}=10^{-4}$. However, for larger values of $\mathcal{P}^{\rm max}_{\zeta,\rm tree}$, the inflaton mean velocity $\langle\dot{\phi}\rangle$ in the lattice simulation deviates from SPT predictions, with O(1) deviations for $\mathcal{P}^{\rm max}_{\zeta,\rm tree}=1$. In particular, during the USR phase, the velocity in the simulations does not decrease as much as predicted by the purely homogeneous evolution. This is a consequence of the backreaction of fluctuations on the background evolution, which is not taken into account by SPT. In Sec. III.2.3, we give tentative explanations of this phenomenon. Due to the very small inflaton velocity throughout the USR phase, the backreaction effect in the mean field value $\langle\phi\rangle$ as a function of number of $e$-folds is much less significant, but the two are naturally related, as the phase-space portrait in Fig. 10 will show explicitly. In particular, the field evolution returns to the SR attractor after the USR phase.
@@ IV Conclusions and outlook
It was recently suggested that USR dynamics advocated by some of the PBH formation scenarios may violate perturbativity and induce loop corrections affecting much longer modes associated to the scales observed through the CMB Kristiano and Yokoyama (2024c). [^7] Instead, see Fumagalli et al. (2024) for effects at near infrared scales when the power spectrum enhancement is due to particle production. While the existence and magnitude of this effect is still under debate Cheng et al. (2022); Riotto (2023a); Kristiano and Yokoyama (2024d); Riotto (2023b); Kristiano and Yokoyama (2024b); Franciolini et al. (2024b); Davies et al. (2024); Firouzjahi and Riotto (2024); Iacconi et al. (2024); Motohashi and Tada (2023); Tasinato (2024, 2023); Firouzjahi (2023); Fumagalli (2023); Cheng et al. (2024); Tada et al. (2024); Firouzjahi (2024); Braglia and Pinol (2024); Kawaguchi et al. (2024a); Ballesteros and Egea (2024); Inomata (2024); Fumagalli (2024); Kawaguchi et al. (2024b); Green and Gupta (2024), we are not able to address this question due to the relatively small range of scales that can be tracked in the lattice box. However, we are in the position to comment on some of the related issues. In particular, our result shows perturbativity is still not completely violated as long as the amplitude of perturbations are smaller than order unity (contrary to what happens in oscillatory potentials, where perturbativity is completely broken already for $\mathcal{P}_{\zeta,\rm tree}^{\rm max}=10^{-2}$ Caravano et al. (2024)). Also, we confirm the existence of important backreaction on the background dynamics, typically captured by tadpole contributions, which were suggested to play a crucial role in protecting long modes from loop corrections Inomata (2024); Fumagalli (2024).
@@
Looking forward, it would be interesting to extend the present work along different directions. Firstly, it will be important to derive the full non-Gaussian statistics for the curvature perturbation, including the nonlinear relation between $\zeta$ and $\delta\phi$ from our lattice simulations, to assess whether positive NGs in $\zeta$ are obtained in all scenarios considered here, as suggested in Ref. Firouzjahi and Riotto (2023). Other groups attempted to study the effect of backreaction and nonlinearities in the USR scenarios employing full in-in computations, $\delta N$, or the stochastic inflation formalism (see e.g. Pattison et al. (2017); Biagetti et al. (2018); Ezquiaga et al. (2020); Ballesteros et al. (2020); Pattison et al. (2021); Cai et al. (2022); Biagetti et al. (2021); Figueroa et al. (2022); Raatikainen et al. (2024); Mizuguchi et al. (2024); Ballesteros and Egea (2024); Animali and Vennin (2024); Jackson et al. (2024)), and it would be interesting to compare these techniques with lattice results. Additionally, evaluating the four-point function of the field fluctuations $\delta\phi$ would provide key insights into the non-perturbative source of SIGW (complementing recent efforts in this direction Cai et al. (2019); Unal (2019); Yuan and Huang (2021); Atal and Domènech (2021); Adshead et al. (2021); Abe et al. (2023); Chang et al. (2023); Garcia-Saenz et al. (2023); Li et al. (2023); Perna et al. (2024); Ruiz and Rey (2024)). These next steps will allow us to better quantify the importance of nonlinearities in the USR scenarios, paving the way for robust observational predictions for GW experiments, as well as constraints on the PBH abundance.
@@ Acknowledgements.
We especially thank Sebastian Zell and Guillermo Ballesteros for proposing to simulate USR prior to this work and for interesting discussions. We also thank Keisuke Inomata, Yuichiro Tada and Vincent Vennin for insightful discussions. A.C. acknowledges funding support from the Initiative Physique des Infinis (IPI), a research training program of the Idex SUPER at Sorbonne Université. During the course of this work, S.RP was supported by the European Research Council under the European Union’s Horizon 2020 research and innovation programme (grant agreement No 758792, Starting Grant project GEODESI). This article is distributed under the Creative Commons Attribution International Licence (CC-BY 4.0).